# Quantum anomalous Hall phases in gated rhombohedral graphene

### v1

@@ Abstract
We consider a coupled system of Dirac operators that finds applications as a macroscopic model of spin and valley polarized gated rhombohedral graphene (RHG) with an arbitrary number of layers as well as in replica models of Floquet topological insulators. We classify all quantum anomalous Hall phases that are compatible with the model and show that a bulk-edge correspondence between bulk phases and chiral edge states carrying a quantized anomalous Hall charge applies. When the displacement field is sufficiently small compared to the interlayer coupling in the RHG application, we retrieve the known phases where the charge is given by the number of graphene layers. When the displacement field increases, we identify all possible topological phase transitions and corresponding quantized chiral edge charges. Numerical simulations confirm the theoretical findings.
Keywords: Topological insulators, graphene systems, Quantum anomalous Hall effect
@@ 1 Introduction
Since the discovery of the quantum Hall effect [20, 30, 1, 9, 25], it was recognized that many phases of matter in different fields of physics were topological in origin [10, 14, 16, 29, 21]. In topological insulators, a distinct manifestation of the topological nature of the phases is a transport asymmetry at interfaces separating insulators in different topological phases. This asymmetry is typically related to the bulk topological phases by a bulk edge correspondence (BEC) [9, 25]. We focus here on effective, macroscopic models of two dimensional systems. In that context, the BEC was demonstrated to hold for a class of effective elliptic problems including those considered in this paper in [3, 26, 4, 5]; see [11, 29, 25] for derivations of bulk-edge correspondences in different settings. A salient feature of the work [3, 26] is the observation of that bulk phase differences are more generally defined than differences of (possibly ill-defined) bulk phases; see also [2, 28, 29] for related works where bulk phases may be defined. The physical transport asymmetry is then related to a topological bulk-difference invariant (BDI), which remains to be computed for problems of interest.
@@
One such class of topological phases is referred to as the quantum anomalous Hall effect (QAHE), first proposed in a prototype model in [15] and analyzed for effective models of multi-layered systems in a series of works; see, e.g., [23, 32, 31]. The first experimental result observing this quantized Hall conductance was obtained in [17] for a rhombohedral graphene (RHG) system. This system, assumed to be spin-polarized and valley-polarized, forms the main application for the effective model described in section 2.
@@
A second application may be found in the field of Floquet topological insulators (FTI). When a sheet of graphene is irradiated by a time-harmonic laser field, effective replica models may be obtained asymptotically in inverse powers of the laser frequency. These models have the same mathematical form as those in RHG applications. See [24, 22] for details on this application and [7] for models of the form (1) considered below. For concreteness, we focus on the RHG application in this paper.
@@
The effective Hamiltonian we consider in this paper models an arbitrary number of layers in gated RHG. It includes two main parameters, the interlayer coupling coefficient, and the difference of potential across the layers generated by a constant displacement field. The objective of this paper is a complete classification of the QAHE phases as these parameters vary. For small values of the displacement field compared to the coupling constant, we retrieve a transport asymmetry given by the number of layers in the system. For large values of the displacement field, we retrieve results obtained in [7] in the context of FTI.
The classification is based on explicit computations of the BDI. This is obtained for systems with rotational symmetries, which highly simplify the computation of Chern numbers as in [29, 12]. This uses the standard result that eigenvalues of Jacobi matrices (tridiagonal symmetric matrices with positive off-diagonal coefficients) are simple. Numerical simulations of interface Hamiltonians provide a quantitative description of the edge modes for different values of the number of layers and the displacement field.
@@
An outline for the rest of the paper is as follows. Section 2 presents the model Hamiltonians and the main results on their topological classification. The derivation of these results is postponed to section 3. The numerical scheme used to diagonalize the interface Hamiltonian is described in section 4. Concluding remarks are given in section 5 while relevant results on the classification of elliptic Hamiltonians are recalled in Appendix A.
@@ 2 Effective model and classification
where $D\cdot\sigma=-i\partial_{x}\sigma_{1}-i\tau\partial_{y}\sigma_{2}$ for $(x,y)\in\mathbb{R}^{2}$ parametrizing the two-dimensional material, $\sigma_{1,2}$ are standard Pauli matrices, and $\tau=\pm 1$ is a valley index. For the rest of the paper, we assume the Fermi velocity $v_{0}$ constant and normalized to 11 by an appropriate rescaling of the spatial variables. We also assume a valley index $\tau=1$ unless otherwise mentioned knowing that all invariants presented below should be multiplied by ${\rm sign}(\tau)$. The real numbers $u_{j}$ for $1\leq j\leq m$ describe the electric potential (bias) at layer $j$.
@@
with $\gamma$ a parameter of the model, which in [17, SM] has approximate value $\gamma=0.435\,$eV. Here, denotes Hermitian transpose. The above choices of coupling correspond to AB and BA stacking, respectively.
@@
We note that $u_{m+1-j}=-u_{j}$ for $1\leq j\leq m$, which will allow to show that the model satisfies the particle-hole symmetry (symmetry of the spectrum about 0).
Note also that the Hamiltonian $H$ is tridiagonal when $B=\gamma A$, which we assume from now on unless mentioned otherwise.
@@
The bulk Hamiltonians $H_{B}$ are defined as the model $H$ of (1) with all coefficients constant. We denote the two bulk operators $H_{B}^{N/S}$ corresponding to $u^{N}=u>0$ for $H_{B}^{N}$ and $u^{S}=-u$ for $H_{B}^{S}$ while all coefficients $u_{j}$ are given by (3). ($N/S$ standing for North/South.)
The Interface Hamiltonian $H_{I}$ is defined as $H$ in (1) with now $u_{j}=u_{j}(y)$. We model a transition from one given displacement field generating the potential $u(y)=u$ for $y\geq R>0$ to an opposite displacement field generating u(y)=uu(y)=-u for $y\leq-R$.
@@
Our objective is to propose a classification of $H_{I}$ based on the transport asymmetry observed along the interface $y\approx 0$. This quantized asymmetry is independent of the profile $u(y)$ so long as $u>0$.
@@
$H_{B}=\mathcal{F}^{-1}\hat{H}_{B}\mathcal{F}$ (4)
where $\mathcal{F}$ is Fourier transform and $\hat{H}_{B}({\mathbf{k}})$ is a $2m\times 2m$ Hermitian matrix where $D_{x}$ and $D_{y}$ are replaced by $k_{x}$ and $k_{y}$ with ${\mathbf{k}}=k_{x}+ik_{y}\in\mathbb{C}$ (identified with $(k_{x},k_{y})\in\mathbb{R}^{2}$ whenever necessary). The spectrum of $H_{B}$ is then given by $2m$ branches of (absolutely continuous) spectrum $\mathbf{k}\mapsto E_{j}(\mathbf{k})$ parametrized by $\mathbf{k}$$\in\mathbb{C}$. These branches are in fact analytic as we will show that they are simple [19, Theorem VII.1.7 and Section VII.3.1].
@@
Define the $2m\times 2m$ matrices $\Gamma_{j}={\rm Diag}(\sigma_{j})$ for $j=1,2,3$ and $G_{1}$ the matrix with $\sigma_{1}$ on the (block) antidiagonal. We denote by $\hat{H}_{B}({\mathbf{k}},\gamma,u)$ the operator constructed with parameters $(\gamma,u)$ and $\hat{H}_{B}(k,\gamma,u)$ to be $\hat{H}_{B}({\mathbf{k}},\gamma,u)$ with $\theta=0$ so that $\mathbf{k}=|\mathbf{k}|=k\in\mathbb{R}$.
@@
$U^{*}(\theta)\hat{H}({\mathbf{k}})U(\theta)$ $=$ $\hat{H}(k),\quad G_{1}\hat{H}(k,u)G_{1}=\hat{H}(k,-u),$ (6)
$G_{1}\Gamma_{3}(-\hat{H}(k))\Gamma_{3}G_{1}$ $=$ $\hat{H}(k),\quad\Gamma_{3}\hat{H}(k,\gamma)\Gamma_{3}=H(-k,-\gamma).$ (7)
The proof of this result is a simple verification using that $H(k,\gamma,u)=H(-k,-\gamma,-u)$. This shows that $\hat{H}(\mathbf{k})$ and $\hat{H}(\mathbf{k})$ are unitarily equivalent implying the parity relation $E_{2m+1-j}({\mathbf{k}})=-E_{j}({\mathbf{k}})$. These relations also show that the branches of spectrum of $H_{B}$ are invariant by rotation.
@@
The above expressions were written for $\tau=1$ and $B=\gamma A$. For ${\mathbf{k}}=ke^{i\theta}$, we observe that
$\hat{H}(\tau=-1,{\mathbf{k}})$ $=$ $\hat{H}(\tau=1,{\mathbf{k}}^{*})=U^{*}(\theta)\hat{H}(\tau=1,k)U(\theta)$ (8)
$\hat{H}(\gamma A^{*},{\mathbf{k}})$ $=$ $\Gamma_{1}\hat{H}(\gamma A,{\mathbf{k}}^{*})\Gamma_{1}=\Gamma_{1}U^{*}(\theta)\hat{H}(\gamma A,k)U(\theta)\Gamma_{1}.$
The spectrum of $H$ is therefore invariant with respect to the map $A\to A^{*}$ (corresponding to a BA stacking rather than an AB stacking). The unitary transformation from $\hat{H}({\mathbf{k}})$ to $\hat{H}(k)$ encoded by $U(\theta)$ in (5) is therefore replaced by $U^{*}(\theta)$ when $\tau=-1$ and $B=\gamma A$ and by $\Gamma_{1}U^{*}(\theta)$ when $\tau=1$ and $B=\gamma A^{*}$. Conjugation by $\Gamma_{1}U(\theta)$ similarly maps $\hat{H}(\tau=-1,\gamma A^{*},{\mathbf{k}})$ to $\hat{H}(\tau=-1,\gamma A^{*},k)$.
@@
Let $\hat{H}_{B}^{h}({\mathbf{k}})$ be defined as above for $\gamma>0$ and $u>0$ with $h\in\{N,S\}$. Let $k=|{\mathbf{k}}|>0$. Then there is $0<E_{0}=E_{0}(\gamma,u,k)$ such that $\hat{H}_{N}^{h}$ is gapped in $(-E_{0},E_{0})$. Moreover, all eigenvalues of the operator $\hat{H}^{h}({\mathbf{k}})$ are simple.
Proof. By rotational invariance we may assume that ${\mathbf{k}}=k$ for $k>0$. Then, $\hat{H}(k)$ is a symmetric (Jacobi) tridiagonal matrix with positive off-diagonal entries given by $k>0$ and $\gamma>0$. We know that their eigenvalues are simple (since the first component of any eigenvector uniquely determines the other components iteratively). Since $E_{2m+1-j}({\mathbf{k}})=-E_{j}({\mathbf{k}})$ and $|E_{j}|({\mathbf{k}})\to\infty$ as $|{\mathbf{k}}|\to\infty$ (by ellipticity of $H_{B}$), we deduce that $E_{j}({\mathbf{k}})=0$ is not possible unless ${\mathbf{k}}=0$ and hence the existence of a gap $E_{0}(\gamma,u,k)>0$.
This result shows that gap closing at $E=0$ can only occur when $k=0$. Such closings do occur and induce (bulk and edge) phase transitions. When no such gap closing occurs at ${\mathbf{k}}=0$, then the spectral gap is global in the sense that there exists $0<E_{0}(\gamma,u)$ such that the spectrum $H_{B}$ does not intersect $(-E_{0},E_{0})$ (by a compactness argument since $|E_{j}({\mathbf{k}})|\to\infty$ as $|{\mathbf{k}}|\to\infty$ and ${\mathbf{k}}\to E_{j}({\mathbf{k}})$ is continuous; in fact real-analytic [19, Theorem VII.1.7 and Section VII.3.1]).
For such values of $(u,\gamma)$, we therefore unambiguously define the projectors
@@
for $h\in\{N,S\}$ projecting onto the $m$- dimensional vector space spanned by the eigenvectors associated to the negative eigenvalues of $\hat{H}^{h}_{B}({\mathbf{k}})$.
We next define the bulk-difference invariant (BDI) [3]
${\rm BDI}(\gamma,u)={\mathfrak{C}}[\Pi^{S},\Pi^{N}]=\frac{i}{2\pi}\int_{\mathbb{R}^{2}}{\rm tr}\Pi^{S}d\Pi^{S}\wedge d\Pi^{S}-\frac{i}{2\pi}\int_{\mathbb{R}^{2}}{\rm tr}\Pi^{N}d\Pi^{N}\wedge d\Pi^{N}.$ (10)
This invariant may be interpreted as the Chern number of the family of projectors $\{\Pi^{N},\Pi^{S}\}$ defined on a unit sphere and is thus guaranteed to take values in the integers. For the operators considered in this paper, the integrals of the Berry curvature $\Pi^{h}d\Pi^{h}\wedge d\Pi^{h}$ appropriately normalized are not guaranteed to be integer-valued [2]. Whereas absolute phases for the N/S insulators may thus not be defined unambiguously, phases differences as in (10) are indeed well-defined; see [7, 5] and the appendix for more detail.
The Hamiltonian $H_{I}$ and corresponding bulk insulators $H^{N/S}_{B}$ satisfy the ellipticity conditions of [3, 26]. As a consequence, the bulk-edge correspondence [3, 26] applies and the asymmetry of $H_{I}$ (heuristically defined as the number of edges modes propagating in one direction in excess of those propagating in the opposite direction) is then exactly given by the above BDI.
We refer to the Appendix for a summary of results leading to the definition of (10), of the edge invariant $2\pi\sigma_{I}[H_{I}]$, and the bulk-edge correspondence stating that $2\pi\sigma_{I}[H_{I}]={\rm BDI}(\gamma,u)$.
@@
It remains to compute the BDI as a function of $(\gamma,u)$ provided that $E_{0}(\gamma,u)>0$ and to identify the values of $(\gamma,u)$ where phase transitions may occur. The number of possible such phases depends on the number of layers $m$.
At $k=0$, the matrix $\hat{H}(\mathbf{k})$ split into $1\times 1$ or $2\times 2$ blocks. The latter are of the form $(u_{j},\gamma;\gamma,u_{j+1})$ for $1\leq j\leq m-1$. The eigenvalues of these blocks cross $E=0$ when $u_{j}u_{j+1}=\gamma^{2}$. Assuming a constant displacement field, this is the constraint
@@
Each eigenvalue $E=0$ is degenerate of multiplicity $2$. Since $\gamma>0$ implies $\delta>\frac{1}{2}$, we deduce that $\delta$ takes the values $j-\frac{m}{2}$ for $\lceil\frac{m}{2}+1\rceil\leq j\leq m-1$. This is $\frac{1}{2}(m-3)$ values when $m$ is odd and $\frac{m}{2}-1$ values when $m$ is even, or $\left\lfloor\frac{m}{2}-1\right\rfloor$ in both cases. We thus observe a first transition at $\delta=1$ when $m=4$ and $\delta=\frac{3}{2}$ when $m=5$.
We define

### v2

@@ Abstract
We consider a coupled system of Dirac operators that models spin- and valley-polarized gated rhombohedral graphene (RHG) with an arbitrary number of layers. We classify all quantum anomalous Hall phases that are compatible with the model and show that a bulk-edge correspondence exists between bulk phases and chiral edge states carrying a quantized anomalous Hall current. When the displacement field is sufficiently small compared to the interlayer coupling in the RHG application, we retrieve the known phases where the charge is given by the number of graphene layers. When the displacement field increases, we identify all possible topological phase transitions and corresponding quantized chiral edge charges. Numerical simulations confirm the theoretical findings.
@@ 1 Introduction
Since the discovery of the quantum Hall effect [24, 35, 1, 9, 29], it was recognized that many phases of matter in different fields of physics were topological in origin [14, 18, 19, 33, 25]. In topological insulators, a distinct manifestation of the topological nature of the phases is a transport asymmetry at interfaces separating insulators in different topological phases. This asymmetry is typically related to the bulk topological phases by a bulk edge correspondence (BEC) [9, 29]. We focus here on effective, macroscopic models of two dimensional systems. In that context, the BEC was demonstrated to hold for a class of effective elliptic problems, including those considered in this paper, in Refs [6, 30, 7, 8]; see Refs [15, 33, 29] for derivations of bulk-edge correspondences in different settings. A salient feature of the work in [6, 30] is the observation that bulk phase differences are more generally defined than differences of (possibly ill-defined) bulk phases; see also [5, 32, 33] for related works where bulk phases may be defined. The physical transport asymmetry is then related to a topological bulk-difference invariant (BDI), which remains to be computed for problems of interest.
@@
One such class of topological phases produce quantized electron transport referred to as the quantum anomalous Hall effect (QAHE). This theory, first proposed in a prototype model by Haldane[20], describes quantized electron transport in 2d materials which, unlike in the quantum Hall effect, does not require an external magnetic field to be applied. A number of mechanisms have been used to realize the QAHE [10], and one recent realization is in layered graphene systems with ABC-type, or rhombohedral, stacking. The band structure for effective models of multi-layered graphene systems was analyzed in a series of works [27, 38, 36, 37] and a first experimental result observing quantized Hall conductance in multi-layered graphene at zero magnetic field was obtained in [21] for a five layer rhombohedral graphene (RHG) system and [11] for a 4-layer system. This system, assumed to be spin-polarized and valley-polarized, forms the main application for the effective model described in section 2.
@@
A second application may be found in the field of Floquet topological insulators (FTI). When a sheet of graphene is irradiated by a time-harmonic laser field, effective replica models may be obtained asymptotically in inverse powers of the laser frequency. These models have the same mathematical form as those in RHG applications. See [28, 26] for details on this application and [3] for models of the form (1) considered below. For concreteness, we focus on the RHG application in this paper, though pertinent results from FTI’s are cited where they coincide with the RHG model.
@@
The effective Hamiltonian we consider in this paper models an arbitrary number of layers in gated RHG. It includes two main parameters, the interlayer coupling coefficient, and the difference of potential across the layers generated by a constant displacement field. The objective of this paper is a complete classification of the QAH phases as these parameters vary. For small values of the displacement field compared to the coupling constant, we retrieve a transport asymmetry given by the number of layers in the system. For large values of the displacement field, we retrieve results obtained in [3] in the context of FTI.
As mentioned above, we use the notion of BDI’s to define Chern numbers which topologically characterize the shared band gap of two bulk QAH phases, defined by different gate potentials or inter-layer couplings, which meet along an interface. We prefer this approach to the prevailing approach of defining band-by-band bulk invariants as integrals of Berry curvature for two important reasons. First, even as Haldane originally noted[20], these bulk invariants are not gauge-invariant and thus predictions of absolute values of Hall conductance cannot only rely on the assignment of bulk invariants but must rely on other model considerations to fix a gauge. Second, integrals of Berry curvature are not guaranteed to be bona fide Chern numbers unless the underlying connection is continuous (modulo a smooth gauge transformation) at every point on a closed manifold. For periodic models with a compact Brillouin zone these criteria are usually self-evident, however for continuum models we must consider a non-trivial compactification of the entire plane $\mathbb{R}^{2}\;$[32, 7, 6]. Examples from equatorial waves [34], continuum photonics[32], and the 2d Dirac model itself [5] show that bulk invariants for continuous models often do not define bona fide Chern numbers without regularizing terms, and indeed this is also the case for the RHG model. By considering BDI’s instead of bulk invariants we eliminate the need to fix the correct gauge by considering differences in Berry curvatures and considerably simplify the construction of well-defined Chern numbers by providing much more flexible criterion for their definition (see Appendix A and [6]).
@@
An outline for the rest of the paper is as follows. Section 2 presents the model Hamiltonians and the main results on their topological classification. We highlight the symmetries of the system, which greatly simplify the computation of Chern numbers [33, 16], while noting where simplifications in our model may provide obstructions to the observation of QAH phases in real systems. The derivation of these results is postponed to section 3, where we present a proof of our main theorem defining a BEC for the system. In section 4 we detail a numerical scheme used to approximate the spectrum of continuous Hamiltonians and present numerical simulations of the RHG interface Hamiltonians to provide a quantitative description of the edge modes for different values of the number of layers and the displacement field. Concluding remarks are given in section 5 while relevant results on the classification of elliptic Hamiltonians are recalled in Appendix A.
@@ 2 Effective model and classification
which acts on state vectors $\psi\in H^{1}(\mathbb{R}^{2};\mathbb{C}^{2m})$ for which the coordinates $(\psi_{2j-1},\psi_{2j}),j\in\{1,...,m\}$ describe a pseudo-spin coordinate for, respectively, the $A$ and $B$ sub-lattices of the $j$-th layer[38, 36]. Here $D\cdot\sigma=-i\partial_{x}\sigma_{1}-i\tau\partial_{y}\sigma_{2}$ for $(x,y)\in\mathbb{R}^{2}$ parametrizing the two-dimensional material, $\sigma_{1,2}$ are standard Pauli matrices, and $\tau=\pm 1$ is a valley index. For the rest of the paper, we assume the Fermi velocity $v_{0}$ constant and normalized to 11 by an appropriate rescaling of the spatial variables. We also assume a valley index $\tau=1$ unless otherwise mentioned knowing that all invariants presented below should be multiplied by ${\rm sign}(\tau)$. The real numbers $u_{j}$ for $1\leq j\leq m$ describe the electric potential (bias) at layer $j$.
@@
with $\gamma$ a parameter of the model, which in [21](SM) has approximate value $\gamma=0.435\,$eV. Here, denotes Hermitian transpose. The above choices of coupling correspond to ABC and CBA stacking, respectively- both rhombohedral stacking but with oppostite chirality. See [27] for a detailed discussion of graphene stacking orders.
@@
We note that $u_{m+1-j}=-u_{j}$ for $1\leq j\leq m$, which will allow us to show that the model satisfies particle-hole symmetry. Note also that the Hamiltonian $H$ is tridiagonal when $B=\gamma A$, which we assume from now on unless mentioned otherwise.
@@
The bulk Hamiltonians $H_{B}$ are defined as the model $H$ of (1) with all coefficients constant. We denote the two bulk operators $H_{B}^{N/S}$ corresponding to a gate potential $u=u^{N}$ for $H_{B}^{N}$ and $u=u^{S}$ for $H_{B}^{S}$ while all coefficients $u_{j}$ are given by (3) ($N/S$ standing for North/South). We consider $\gamma$ to be a given constant throughout the rest of the paper.
The interface Hamiltonian $H_{I}$ is defined as $H$ in (1) with now $u_{j}=u_{j}(y)$. We model a transition from one given displacement field generating the potential $u(y)=u^{N}$ for $y\geq R>0$ to another displacement field generating $u(y)=u^{S}$ for $y\leq-R$, assuming that $u(y)$ is smooth in the interval $y\in[-R,R]$.
@@
Our objective is to propose a classification of $H_{I}$ based on the transport asymmetry observed along the interface $y\approx 0$. This quantized asymmetry is independent of the profile u(y)u(y)for any given values of $u^{N},u^{S}$.
@@
$H_{B}=\mathcal{F}^{-1}\hat{H}_{B}({\mathbf{k}})\mathcal{F}$ (4)
where $\mathcal{F}$ is Fourier transform and $\hat{H}_{B}({\mathbf{k}})$ is a $2m\times 2m$ Hermitian matrix where $D_{x}$ and $D_{y}$ are replaced by $k_{x}$ and $k_{y}$ with ${\mathbf{k}}=k_{x}+ik_{y}\in\mathbb{C}$ (identified with $(k_{x},k_{y})\in\mathbb{R}^{2}$ whenever necessary). The spectrum of $H_{B}$ is then given by $2m$ branches of (absolutely continuous) spectrum $\mathbf{k}\mapsto E_{j}(\mathbf{k})$ parametrized by $\mathbf{k}$$\in\mathbb{C}$. These branches are in fact analytic as we will show that they are simple (see [23, Theorem VII.1.7 and Section VII.3.1]).
@@
Define the $2m\times 2m$ matrices $\Gamma_{j}={\rm Diag}(\sigma_{j})$ for $j=1,2,3$ and $G_{1}$ the matrix with $\sigma_{1}$ on the (block) antidiagonal. We denote by $\hat{H}_{B}({\mathbf{k}},\gamma,u)$ the operator $\hat{H}_{B}(\mathbf{k})$ constructed with parameters $(\gamma,u)$ and $\hat{H}_{B}(k,\gamma,u)$ to be $\hat{H}_{B}({\mathbf{k}},\gamma,u)$ with $\theta=0$ so that $\mathbf{k}=|\mathbf{k}|=k\in\mathbb{R}$.
@@
$U^{*}(\theta)\hat{H}({\mathbf{k}},\gamma,u)U(\theta)$ $=$ $\hat{H}(k,\gamma,u),\quad G_{1}\hat{H}(k,\gamma,u)G_{1}=\hat{H}(k,\gamma,-u),$ (6)
$G_{1}\Gamma_{3}(-\hat{H}(k,\gamma,u))\Gamma_{3}G_{1}$ $=$ $\hat{H}(k,\gamma,u),\quad\Gamma_{3}\hat{H}(k,\gamma,u)\Gamma_{3}=H(-k,-\gamma,u).$ (7)
The proof of this result is a simple verification using that $\hat{H}_{B}(k,\gamma,u)=\hat{H}_{B}(-k,-\gamma,-u)$. This shows that $\hat{H}_{B}(\mathbf{k},\gamma,u)$ and $\hat{H}_{B}(\mathbf{k},\gamma,u)$ are unitarily equivalent implying the parity relation $E_{2m+1-j}({\mathbf{k}})=-E_{j}({\mathbf{k}})$. These relations also show that the branches of spectrum of $H_{B}$ are invariant by rotation.
@@
Although we assume for most of our analysis that $\tau=1$ and $B=\gamma A$ we also note that
$\hat{H}_{B}({\mathbf{k}},\tau=-1)$ $=$ $\hat{H}_{B}({\mathbf{k}}^{*},\tau=1)=U^{*}(\theta)\hat{H}_{B}(k,\tau=1)U(\theta)$ (8)
$\hat{H}_{B}({\mathbf{k}},B=\gamma A^{*})$ $=$ $\Gamma_{1}\hat{H}_{B}({\mathbf{k}}^{*},B=\gamma A)\Gamma_{1}=\Gamma_{1}U^{*}(\theta)\hat{H}_{B}(k,B=\gamma A)U(\theta)\Gamma_{1}.$
The spectrum of $\hat{H}_{B}$ is therefore invariant with respect to the map $A\to A^{*}$ (corresponding to a CBA stacking rather than an ABC stacking). The unitary transformation from $\hat{H}_{B}({\mathbf{k}})$ to $\hat{H}_{B}(k)$ encoded by $U(\theta)$ in (5) is therefore replaced by $U^{*}(\theta)$ when $\tau=-1$ and $B=\gamma A$ and by $\Gamma_{1}U^{*}(\theta)$ when $\tau=1$ and $B=\gamma A^{*}$. Conjugation by $\Gamma_{1}U(\theta)$ similarly maps $\hat{H}({\mathbf{k}},\tau=-1,\gamma A^{*})$ to $\hat{H}(k,\tau=-1,\gamma A^{*})$. These unitary equivalences will allow us to extend the BEC to both stacking orders and valley indices without additional analysis.
@@
Let $\hat{H}_{B}^{h}({\mathbf{k}})$ be defined as above for $\gamma>0$ and $|u|>0$ with $h\in\{N,S\}$. Let $k=|{\mathbf{k}}|>0$. Then there is $0<E_{0}=E_{0}(k,\gamma,u)$ such that $\hat{H}_{B}^{h}({\mathbf{k}})$ is gapped in $(-E_{0},E_{0})$. Moreover, all eigenvalues of the operator $\hat{H}_{B}^{h}({\mathbf{k}})$ are simple.
Proof. By rotational invariance we may assume that ${\mathbf{k}}=k$ for $k>0$. Then, $\hat{H}(k)$ is a symmetric (Jacobi) tridiagonal matrix with positive off-diagonal entries given by $k>0$ and $\gamma>0$. We know that their eigenvalues are simple (since the first component of any eigenvector uniquely determines the other components iteratively). Since $E_{2m+1-j}({\mathbf{k}})=-E_{j}({\mathbf{k}})$ and $|E_{j}({\mathbf{k}})|\to\infty$ as $|{\mathbf{k}}|\to\infty$ (by ellipticity of $H_{B}$), we deduce that $E_{j}({\mathbf{k}})=0$ is not possible unless ${\mathbf{k}}=0$ and hence the existence of a gap $E_{0}(k,\gamma,u)>0$.
This result shows that gap closing at $E=0$ can only occur when $\mathbf{k}=0$. Such closings do occur and induce (bulk and edge) phase transitions. When no such gap closing occurs at ${\mathbf{k}}=0$, then the spectral gap is global in the sense that there exists $0<E_{0}(\gamma,u)$ such that the spectrum $\hat{H}^{h}_{B}({\mathbf{k}})$ does not intersect $(-E_{0},E_{0})$ for any $\mathbf{k}$} (by a compactness argument since $|E_{j}({\mathbf{k}})|\to\infty$ as $|{\mathbf{k}}|\to\infty$ and ${\mathbf{k}}\to E_{j}({\mathbf{k}})$ is continuous; in fact real-analytic; (see [23, Theorem VII.1.7 and Section VII.3.1])).
For such values of $(u^{h},\gamma)$, we therefore unambiguously define the projectors
@@
for $h\in\{N,S\}$ projecting onto the $m$-dimensional vector space spanned by the eigenvectors associated to the negative eigenvalues of $\hat{H}^{h}_{B}({\mathbf{k}})$. Recall that we assume $\gamma$ does not vary, while $\hat{H}_{B}^{N}(\mathbf{k})$ is characterized by a gate potential $u^{N}$ and $\hat{H}_{B}^{S}(\mathbf{k})$ is characterized by a gate potential $u^{S}$.
We define the bulk-difference invariant (BDI) [6] as follows:
${\rm BDI}(\gamma,u^{S},u^{N})={\mathfrak{C}}[\Pi^{S},\Pi^{N}]:=\frac{i}{2\pi}\int_{\mathbb{R}^{2}}{\rm tr}\Pi^{S}d\Pi^{S}\wedge d\Pi^{S}-\frac{i}{2\pi}\int_{\mathbb{R}^{2}}{\rm tr}\Pi^{N}d\Pi^{N}\wedge d\Pi^{N}.$ (10)
Although there are no general guarantees that integrals of this form are well-defined, given appropriate assumptions on $\Pi^{N},\Pi^{S}$ (e.g. (27) in the following sections) the invariant may be interpreted as the Chern number of the family of projectors $\{\Pi^{N},\Pi^{S}\}$ defined by stereographic projection of $\mathbb{R}^{2}$ onto the unit sphere and is thus guaranteed to take values in the integers by their interpretation as Atiyah-Singer indices[6, 7]. By contrast, for the operators considered in this paper the integrals of the Berry curvature $\Pi^{h}d\Pi^{h}\wedge d\Pi^{h}$ appropriately normalized are not guaranteed to be integer-valued [5], and even when they are, are not gauge-invariant. Therefore, whereas absolute phases for the N/S insulators may not be defined unambiguously, phase differences as in (10) are indeed well-defined; see Ref [3, 8] and the appendix for more detail.
The Hamiltonian $H_{I}$ and corresponding bulk insulators $H^{N/S}_{B}$ satisfy the ellipticity conditions of [6, 30]. As a consequence, the bulk-edge correspondence [6, 30] applies and the asymmetry of electron transport governed by $H_{I}$ is then exactly given by the above BDI. We refer to the Appendix for a summary of results leading to the definition of (10), of the edge invariant $\sigma_{I}[H_{I}]$ quantifying total electron transport along the interface $y\approx 0$, and the bulk-edge correspondence stating that $2\pi\sigma_{I}[H_{I}]={\rm BDI}(\gamma,u^{S},u^{N})$.
@@
It remains to compute the BDI as a function of $(\gamma,u^{S},u^{N})$ provided that $E_{0}(\gamma,u)>0$ and to identify the values of $(\gamma,u)$ where phase transitions may occur. The number of possible such phases depends on the number of layers $m$.
At $k=0$, the matrix $\hat{H}_{B}(\mathbf{k})$ splits into $1\times 1$ or $2\times 2$ blocks. The latter are of the form $(u_{j},\gamma;\gamma,u_{j+1})$ for $1\leq j\leq m-1$. The eigenvalues of these blocks cross $E=0$ when $u_{j}u_{j+1}=\gamma^{2}$. Assuming a constant displacement field, this is the constraint
@@
When this constraint is satisfied each eigenvalue $E=0$ is degenerate of multiplicity $2$ since $u_{m+1-j}=-u_{j}$. Since $\gamma>0$ implies $\delta(\gamma,u)>\frac{1}{2}$, we deduce that $\hat{H}_{B}({\mathbf{k}})$ has a 2-fold degenerate eigenvalue $E=0$ for critical points $u$ for which $\delta(\gamma,u)$ takes the values $j-\frac{m}{2}$ for $\lceil\frac{m}{2}+1\rceil\leq j\leq m-1$. This is $\frac{1}{2}(m-3)$ values when $m$ is odd and $\frac{m}{2}-1$ values when $m$ is even, or $\left\lfloor\frac{m}{2}-1\right\rfloor$ in both cases. Increasing $u>0$ from a small value compared to $\gamma$, we thus observe a first transition at $\delta(\gamma,u)=1$ when $m=4$ and $\delta(\gamma,u)=\frac{3}{2}$ when $m=5$.
Next define